Let X be a random variable with the following probability distribution:

Value x of XP(X=x)
-300.05
-200.20
-100.05
00.15
100.25
200.30

Find the expectation E(X) and variance Var(X) of X. (if necessary, consult a list of formulas.) 
Answer #1

E(X) = \sum X * P(X)

= -30 * 0.25 -20 * 0.20 -10 * 0.05 + 0 * 0.15 + 10 * 0.25 + 20 * 0.30

= 2.5

Var(X) = \sum X2 * P(X) - E(X)2

= [ (-30)2 * 0.25 + (-20)2 * 0.20 + (-10)2 * 0.05 + 0 * 0.15 + 102 * 0.25 + 202 * 0.30 ] - 2.52

= 268.75